OpenStudy (calculusxy):

A triangle that has two sides of equal length is called isosceles. Make up an example of an isosceles triangle, one of whose vertices is (3,5). If you can, find a triangle that does not have any horizontal or vertical sides.

OpenStudy (calculusxy):

@agent0smith

OpenStudy (calculusxy):

I mainly need help with finding a triangle that does not have horizontal or vertical sides. What does that mean?

OpenStudy (calculusxy):

that seems logical

OpenStudy (agent0smith):

 Distance between (3, 5) and (a, 0) must equal distance between (3, 5) and (c, d) for it to be isosceles. Pick a value of \(a\), then you can work on finding \(c\) and \(d\) using distance formula. You'll be able to also pick a reasonable value of \(c\), once you've found the (3, 5) and (a, 0) distance.

OpenStudy (calculusxy):

Could I do something with the pythagorean theorem?

OpenStudy (agent0smith):

Pick a value of a, then find the distance between (3, 5) and (a, 0).

OpenStudy (calculusxy):

ok

OpenStudy (calculusxy):

i chose that a =1 and i had the distance of \(\sqrt{29}\)

OpenStudy (agent0smith):

Now pick a reasonable value of c (it must be somewhat close to x=3, within sqrt29 units, or else the triangle won't be isosceles) Distance between (3, 5) and (c, d) will equal to sqrt 29, solve for d.

OpenStudy (calculusxy):

how did you figure out that the value of c needs to close to x = 3?

OpenStudy (agent0smith):

Draw it and find out. If you make it too far away, will it be isosceles?

OpenStudy (calculusxy):

I am confused. If x = 3, it would be right underneath the point (3,5).

OpenStudy (calculusxy):

I chose x = 5

OpenStudy (agent0smith):

You said you didn't want horizontal or vertical lines, so why would you pick x=3?

OpenStudy (agent0smith):

Distance between (3, 5) and (c, d) will equal to sqrt 29, solve for d.

OpenStudy (calculusxy):

\(\sqrt{29} = \sqrt{(3-5)^2 + (5-d)^2}\)

OpenStudy (calculusxy):

\[(\sqrt{29})^2 = (\sqrt{3-5)^2 + (5-d)^2})^2 \] \(29 = 4 + (5-d)^2\)

OpenStudy (calculusxy):

Then I get that d = 0, 10

OpenStudy (agent0smith):

Yep.

OpenStudy (calculusxy):

But how do I know which one to choose from?

OpenStudy (agent0smith):

Well, they both work to make a triangle...but you didn't want vert/horiz lines... so you should be able to choose based on that.

OpenStudy (calculusxy):

If i have y = 0 then i don't think that it would work because it would have horizontal and vertical sides. I guess 10?

OpenStudy (calculusxy):

would that be correct @agent0smith ?

OpenStudy (agent0smith):

Sure. Draw them and it should look isosceles.

OpenStudy (calculusxy):

Thank you! Can you help me with another question please?

OpenStudy (calculusxy):